The Big Bull investment club has a portfolio that contains only two stocks. At the end of the year, stock A had gone up 8%, and stock B had gone up 10%. The total portfolio, which had a value of $10,000 at the beginning of the year, went up 9.2%. How much is the investment in stock B?
Answer by jorel1380(3719)

m+n=10000
.08m+.10n=.092(10000)
80m+100n=92(10000)=920000
80m+80n=800000
20n=120000
n=6000
m=4000
$6000 was invested in stock B..